# TimeDiT: General-purpose Diffusion Transformers for Time Series Foundation Model

### v1

@@ 4.2 Time Series Diffusion Transformer
Figure 1 shows the overall framework of TimeDiT. We first establish the $\mathbf{M_{obs}}$ and $\mathbf{x}_{0}^{\text{obs}}$ based on the given input from different distributions with multivariate sequences, missing value and multi-resolution by injecting placeholders to standardize the input shape across different time series, facilitating more efficient and consistent processing. Then, the unified time series mask unit adapts to diverse time series scenarios and builds the $\mathbf{x}_{0}^{\text{con}}$, $\mathbf{M}$ and $\mathbf{x}_{0}^{\text{tar}}$, with shape $\mathbb{R}^{B\times L\times K}$, to help TimeDiT learn robust representations in a self-supervised manner by reconstructing the original sequence through denoising $\mathbf{x}_{T}^{\text{tar}}$. After that, the embedding layer directly treats $\mathbf{x}_{0}^{\text{con}}$ and $\mathbf{x}_{0}^{\text{tar}}$ as without any patching, as the diffusion process is designed to handle multivariate input and operate in a continuous token space. By preserving the integrity of the input time series, TimeDiT ensures that the model can effectively capture and utilize the rich information contained within the data. The TimeDiT block’s attention mechanism is designed to autonomously learn cross-channel and temporal correlations through end-to-end training.
Standardize Pipeline
We introduce placeholders within the input sequences to standardize the input shape across different time series, accounting for varying channel numbers $K$ and sequence length $L$. Specifically, we define the maximum channel number $K_{max}$ such that any input with channel $k<K_{max}$ is padded to have $K^{max}$ channels while any input with more than $K_{max}$ channels are segmented into $\lceil\frac{k}{K_{max}}\rceil$ blocks of inputs where each block has $K_{max}$ channels and undergoes independent processing. This segmentation allows our model to manage high-dimensional data efficiently, reducing computational overhead and maintaining relative positional integrity of the data and consistency across inputs. Additionally, for any input with sequence length less than the designated maximum length $L_{max}$, we pad the sequence in the front to achieve the desired length. This standardization is essential for establishing a uniform input structure that enhances processing efficiency and consistency.
Time Series Mask Unit
@@
We propose a unified time series mask mechanism that includes a variety of masks that seamlessly integrates with the model during self-supervised task agnostic pre-training and task specific fine-tuning to cater to diverse time series scenarios. The time series mask unit generate four types of masks: reconstruction mask, stride mask, block mask, and random mask. Firstly, the task-agnostic pre-training aims to improve the overall time series representation by encouraging the model to learn robust and generalizable features from the input data. Secondly, the task-specific training is designed specifically for the most common downstream tasks, including forecasting and imputation, enabling the model to adapt to the unique requirements of each task.
@@
where $r$ is the mask ratio. In addition, for task-specific training and inference, we allow the user to supply customized imputation masks that could simulate the naturally missing data and multi-resolution cases.
Block mask $\mathbf{M}^{\text{B}}$ can be generated via:
@@
where $l$ is the predicted length. We can randomly select $l$ during pretraining and use the designated prediction length during the finetuning and inference stage for specific experiment settings.
Stride mask $\mathbf{M}^{\text{S}}$, a variant of $\mathbf{M}^{\text{B}}$, is placed intermittently within the series and is defined as follows:
$\mathbf{M}^{\text{S}}(x,n_{\text{blocks}})=\begin{cases}1&\left\lfloor\frac{j}{b}\right\rfloor\bmod 2=0\\ 0&\text{otherwise}\end{cases}$ (6)
where $n_{block}$ is the number of blocks into which the sequence is divided; $b=\left\lceil\frac{L}{n_{\text{blocks}}}\right\rceil$ is the length of each block; $j$ is the index of the sequence. $\mathbf{M}^{\text{S}}$is designed for task-agnostic pretraining further to enhance the model’s representation ability by integrating information across non-contiguous parts of the series.
In addition, reconstruction mask $\mathbf{M}^{\text{Rec}}=0$ is used for tasks including synthetic data generation and anomaly detection, where we can directly generate synthetic data or obtain an anomaly score for each temporal position based on the difference between original and reconstructed series.
For the pretraining stage, we random select one conditional mask type from $\mathbf{M}=\{\mathbf{M}^{\text{R}},\mathbf{M}^{\text{S}},\mathbf{M}^{\text{B}},\mathbf{M}^{\text{Rec}}\}$ for each instance. TimeDiT’s goal is to reconstruct the $\mathbf{x}^{\text{tar}}$, defined as $\mathbf{x}_{0}\times(J-\mathbf{M})$ where $J$ is unit matrix. The target sequence is the masked portion of the original sequence. In the finetuning and inference stage, the choice of mask is tailored to align with the specific requirements of the user. This flexibility allows TimeDiT to apply the most appropriate masking strategy based on the context of the task and application.
Condition Injection
Instead of following  (Peebles and Xie, 2022) to integrate diffusion timestep and label information (texts in out case) through the layer normalization preceeding the attention block, we add the diffusion timestep and texts information to the target noise as these are universal information to the series.
Given that TimeDiT utilizes a transformer-based architecture, a straightforward and intuitive approach is to include conditional information directly as part of the input sequence by concatenation as done in latent diffusion  (Rombach et al., 2022). However, we empirically found out that controlling mean and variance through layer normalization is a stronger form of conditional information injection. We incorporate the partial observations $x_{0}^{c}on$ through adaptive layer normalization to control the scale and shift of the $x_{0}^{tar}$. This design choice is motivated by the face that the scale and shift of the partial observations are more relevant to the target observations. This integration can be expressed as
$\text{AdaLN}(h,c)=c_{scale}\text{LayerNorm}(h)+c_{shift}$ (7)
where $h$ is the hidden state and $c_{\text{scale}}$ and $c_{\text{shift}}$ are the scale and shift parameters derived from the partial observations. We perform temporal attention in the self-attention block to capture the temporal dependency within the input.
Inference
We perform pre-training and finetuning for forecasting and imputation. We exclude the pre-training process from anomaly detection and synthetic generation because the two tasks are very dataset specific and do not necessarily benefit from learning distributions beyond the target dataset. Let $n$ represent the number of samples generated for each prediction, which we set to $n=10$ ($n=30$ for forecasting tasks) in our experimental setup at inference time. We use the median of these $n$ predictions as the final prediction, providing the added benefit of obtaining a confidence interval for TimeDiT ’s predictions. To prevent channel padding from affecting the generated samples, we mask out the invalid channels during sampling at each diffusion timestep so that TimeDiT does not falsely treat information in the non-valid channels as meaningful information. Padding is applied at the beginning of the temporal dimension to ensure that the most relevant information remains at the end, thereby mitigating the effect of padding.
4.3 Physics-Informed TimeDiT
Algorithm 1 Physics Informed TimeDiT through Energy-based Sampling
1:$\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
@@ 4.3 Physics-Informed TimeDiT
3:  $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, else $\mathbf{z}=\mathbf{0}$
4:  $\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z}$
@@
6:for $j=0,1,..,k-1$ do
7:  $\mathbf{x}^{tar}_{j+1}=\mathbf{x}^{tar}_{j}+\epsilon\nabla K(\mathbf{x}_{j}^{tar};\mathbf{x}^{obs})+\alpha\epsilon\nabla\log p(\mathbf{x}_{j}^{tar}|\mathbf{x}^{obs})+\sqrt{2\epsilon}\sigma,\sigma\sim\mathcal{N}(0,1)$
@@
9:return $\mathbf{x}^{tar}_{k}$
Physics principles are fundamental in shaping the evolution of temporal signals observed in real-world phenomena, such as climate patterns and oceanographic data. Therefore, it is essential to integrate physical knowledge into foundational time series models. In this section, we propose a strategy to incorporate physics knowledge as an energy-based prior for TimeDiT during inference, which iteratively refines the reverse diffusion process. By guiding the denoising process during inference with gradients derived from physical laws represented by partial differential equations (PDEs), the integration of this knowledge can significantly enhance the quality of the generated samples.
A generic form of a physical law represented as a PDE that describes the evolution of a continuous temporal signal $\mathbf{x}(\mathbf{u},t)$ over a spatial coordinate $\mathbf{u}$ is given by:
@@
This function reaches its maximum when the predicted time series is perfectly consistent with the physical model, resulting in a residual of 0. Using this metric $K$, physics knowledge can be integrated into a probabilistic time series foundation model $p(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})$ by solving the following optimization problem to obtain a refined model $q(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})$:
$q(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})=\arg\max_{q}\mathbb{E}_{\mathbf{x}^{\text{{tar}}}\sim q}K(\mathbf{x}^{\text{{tar}}};F)-\alpha D_{KL}(q(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})||p(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}}))$ (10)
where the first term represents the aforementioned physics knowledge metric, and the second term controls the divergence between $q(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})$ and $p(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})$. The following theorem provides a closed-form solution to the above optimization problem:
@@
The theorem illustrates that sampling from the Boltzmann distribution defined in Eq.  11, is analogous to incorporating physics knowledge into model edition. In the context of diffusion models, this distribution can be effectively sampled using Langevin dynamics (Stoltz et al., 2010):
@@
In diffusion model, precisely calculate the likelihood $\log p(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})$ is intractable. To tackle this issue, following previous works (Kollovieh et al., 2023), we approximate likelihood with the objective to edit the pre-trained diffusion model:
$\log p(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})=-\mathbb{E}_{\epsilon,t}[||\epsilon_{\theta}(\mathbf{x}^{\text{tar}},t;\mathbf{x}^{\text{{con}}})-\epsilon||^{2}]$ (14)
The approximation presented above constitutes the optimizable component of the evidence lower bound(ELBO). Algorithm 1 summarizes the comprehensive model editing process.

### v2

@@ 3.2 Time Series Diffusion Transformer
Figure 1 shows the overall framework of TimeDiT. Firstly, we establish $\mathbf{M_{obs}}$ and $\mathbf{x}_{0}^{\text{obs}}$ based on inputs with varying shapes, missing values, and multi-resolution data. By injecting placeholders, we identify corresponding positions and standardize input shapes across different time series, enabling more efficient and consistent processing. Then, the unified time series mask unit constructs $\mathbf{M}$ and adapts to diverse scenarios, generating $\mathbf{x}_{0}^{\text{con}}$ and $\mathbf{x}_{0}^{\text{tar}}$ with shape $\mathbb{R}^{B\times L\times K}$, where $B$ is the batch size. This enables TimeDiT to learn robust representations in a self-supervised manner by reconstructing the original sequence through denoising $\mathbf{x}_{T}^{\text{tar}}$. Adopting a "What You See Is What You Get" (WYSIWYG) design philosophy, our model represents tokens as direct, contiguous arrays of inputs. After that, the embedding layer with linear projection maps $\mathbf{x}_{0}^{\text{con}}$ and the noised $\mathbf{x}_{0}^{\text{tar}}$ into a continuous token space without vector quantization (Li et al., 2024), thereby preserving input integrity. To model the per-token probability distribution, the TimeDiT block is designed to autonomously learn cross-channel and temporal correlations through end-to-end training.
Diffusion process. TimeDiT unconditional diffusion process comprises a forward process that gradually adds noise to a data sample $x_{0}\sim q(x)$, transforming it into Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$ as defined by Eq. 13 and a reverse denoising process learned by a neural network (Eq. 14). To guide samples toward regions of high classifier likelihood, a self-conditional component $\mathbf{x}^{\text{con}}_{0}$ is integrated. We can train the denoising model $\bm{\mu}_{\theta}\left(\mathbf{x}_{t}^{\mathrm{tar}},\mathbf{x}_{0}^{\mathrm{con}}\right)$ in Eq.  1 using a weighted mean squared error (MSE) loss, which can be justified as optimizing a weighted variational lower bound on the data log-likelihood:
$L(\mathbf{x}^{\text{con}}_{0})=\sum_{t=1}^{T}\mathbb{E}_{q(\mathbf{x}^{\text{tar}}_{t}|\mathbf{x}^{\text{con}}_{0})}\|\mu(\mathbf{x}^{\text{tar}}_{t},\mathbf{x}^{\text{con}}_{0})-\mu_{\theta}(\mathbf{x}^{\text{tar}}_{t},t|\mathbf{x}^{\text{con}}_{0})\|^{2},$ (2)
where $\mu(\mathbf{x}^{\text{tar}}_{t},\mathbf{x}^{\text{con}}_{0})$ is the mean of the posterior $q(x^{\text{tar}}_{t-1}|\mathbf{x}^{\text{con}}_{0},\mathbf{x}^{\text{tar}}_{t})$.
Transformer-based Condition Injection. TimeDiT employs a transformer-based architecture to process multivariate time series data. We feed the embedding of noised target series $\mathbf{x}^{\text{tar}}_{t}$ (with noise schedule $\beta_{t}\in(0,1)$), and conditional observation $\mathbf{x}^{\text{con}}_{0}$ into the TimeDiT block, where the multi-head attention aims to then learns complex relationships within the data. During the diffusion process, unlike previous approaches (Peebles and Xie, 2022; Lu et al., 2024), we innovatively inject diffusion time information directly into the target noise as these represent universal information across the noised series. For self-conditional information, while a straightforward approach would be to include conditional information directly in the input sequence through concatenation (Rombach et al., 2022), we employ adaptive layer normalization (AdaLN) to control the scale and shift of $x_{0}^{\text{tar}}$ using partial observations $x_{0}^{\text{con}}$:
$\text{AdaLN}(h,c)=c_{scale}\text{LayerNorm}(h)+c_{shift},$ (3)
where $h$ is the hidden state and $c_{\text{scale}}$ and $c_{\text{shift}}$ are the scale and shift parameters derived from the $x_{0}^{\text{con}}$. This method proved empirically more effective than simple input concatenation, as it leverages the scale and shift of $x_{0}^{\text{con}}$, which are crucial for capturing temporal continuity and progression.
@@
Time Series Mask Unit. The Time Series Mask Unit is a key component of our model, designed to enhance its versatility and performance across various time series tasks. This unified mechanism incorporates multiple mask types that seamlessly integrate with the model throughout its lifecycle - from self-supervised task-agnostic pre-training to task-specific fine-tuning and inference. The time series mask unit generates four distinct mask types: random mask $\mathbf{M}^{\text{R}}$, block mask $\mathbf{M}^{\text{B}}$, stride mask $\mathbf{M}^{\text{S}}$, and reconstruction mask $\mathbf{M}^{\text{Rec}}$. During task-agnostic pre-training, these masks help the model develop robust and generalizable features from the input data, improving overall time series representation. In task-specific training, the masks adapt to the unique requirements of common downstream tasks such as forecasting and imputation, enabling the model to specialize effectively.
@@
where $r$ is the mask ratio. For task-specific training and inference, we allow the user to supply customized imputation masks, which replace the random position masks, that could handle the naturally missing data and multi-resolution cases. In addition, block mask $\mathbf{M}^{\text{B}}$ can be generated via:
@@
where $l$ is the predicted length. This mask offers flexibility across different stages of model development and application: during pre-training, a random $l$ exposes the model to various forecasting horizons, while in fine-tuning and inference, a fixed $l$ aligns with specific task requirements. Moreover, stride mask $\mathbf{M}^{\text{S}}$, a variant of $\mathbf{M}^{\text{B}}$, is designed for intermittent placement within time series during task-agnostic pretraining:
$\mathbf{M}^{\text{S}}(x,n_{\text{blocks}})=\begin{cases}1&\left\lfloor\frac{j}{b}\right\rfloor\bmod 2=0\\ 0&\text{otherwise},\end{cases}$ (6)
where $n_{block}$ is the number of blocks; $b=\left\lceil\frac{L}{n_{\text{blocks}}}\right\rceil$ is the length of each block; $j$ is the index of the sequence. It improves the modeling of temporal and inter-correlated dependencies by integrating information across non-contiguous parts of time series, leveraging neighboring values as additional context. In addition, reconstruction mask $\mathbf{M}^{\text{Rec}}=0$ is employed for tasks such as synthetic data generation and anomaly detection. It allows direct generation of synthetic data or calculation of anomaly scores for each temporal position by comparing original and reconstructed series.
3.3 Physics-Informed Sampling
Algorithm 1 Physics-Informed TimeDiT through Energy-based Sampling
1:$\bm{x}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$
@@ 3.3 Physics-Informed Sampling
3:  $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$ if $t>1$, else $\bm{z}=\bm{0}$
4:  $\bm{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\bm{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)\right)+\sigma_{t}\bm{z}$
@@
6:for $j=0,1,\dotsc,k-1$ do
7:  $\bm{x}^{tar}_{j+1}=\bm{x}^{tar}_{j}+\epsilon\nabla K(\bm{x}_{j}^{tar};\bm{x}^{obs})+\alpha\epsilon\nabla\log p(\bm{x}_{j}^{tar}|\bm{x}^{obs})+\sqrt{2\epsilon}\sigma,\sigma\sim\mathcal{N}(0,1)$
@@
9:return $\bm{x}^{tar}_{k}$
Physics principles are fundamental in shaping the evolution of temporal signals observed in real-world phenomena, such as climate patterns and oceanographic data. Therefore, it is essential to integrate physical knowledge into foundational time series models. In this section, we aim at developing a decoding method that can ensure the $\mathbf{x}^{\text{{tar}}}$ generated by TimeDiT to satisfy our prior knowledge to the physical laws. To this end, we propose a strategy to incorporate physics knowledge as an energy-based prior for TimeDiT during inference, which iteratively refines the reverse diffusion process. By guiding the denoising process during inference with gradients derived from physical laws represented by partial differential equations (PDEs), the integration of this knowledge can ensure $\mathbf{x}^{\text{{tar}}}$ to satisfy the PDEs and significantly enhance the quality of the generated samples.
We first start with a brief introduction to physical laws and PDE. A generic form of a physical law represented as a PDE that describes the evolution of a continuous temporal signal $\mathbf{x}(\mathbf{u},t)$ over a spatial coordinate $\mathbf{u}$ is given by:
@@
This function reaches its maximum when the predicted time series is perfectly consistent with the physical model, resulting in a residual of 0. Using this metric $K$, physics knowledge can be integrated into a probabilistic time series foundation model $p(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})$ as an explicit regularization by solving the following optimization problem to obtain a refined model $q(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})$:
$\begin{split}q(\mathbf{x}^{\text{tar}}|\mathbf{x}^{\text{con}})=\arg\max_{q}\Big{[}\mathbb{E}_{\mathbf{x}^{\text{tar}}\sim q}K(\mathbf{x}^{\text{tar}};F)-\alpha D_{KL}(q(\mathbf{x}^{\text{tar}}|\mathbf{x}^{\text{con}})||p(\mathbf{x}^{\text{tar}}|\mathbf{x}^{\text{con}}))\Big{]}\end{split}$ (9)
where the first term represents the aforementioned physics knowledge metric, and the second term controls the divergence between $q(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})$ and $p(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})$. However directly updating the model parameters to optimize the above function is resource-consuming. To solve this issue, we derived the closed-form solution, which does not need updating the model parameters. The above optimization problem has a closed-form solution as provided by the following theorem:
@@
The theorem illustrates that sampling from the Boltzmann distribution defined in Eq.  10, is analogous to incorporating physics knowledge into model edition. In the context of diffusion models, this distribution can be effectively sampled using Langevin dynamics (Stoltz et al., 2010):
@@
where $\sigma\sim\mathcal{N}(0,1)$. In diffusion model, precisely calculate the likelihood $\log p(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})$ is intractable. To tackle this issue, following previous works (Kollovieh et al., 2023), we approximate likelihood with the objective to edit the pre-trained diffusion model: $\log p(\mathbf{x}^{\text{{tar}}}|\mathbf{x}^{\text{{con}}})=-\mathbb{E}_{\epsilon,t}[||\epsilon_{\theta}(\mathbf{x}^{\text{tar}},t;\mathbf{x}^{\text{{con}}})-\epsilon||^{2}]$. The approximation presented above constitutes the optimizable component of the evidence lower bound(ELBO). Algorithm 1 summarizes the comprehensive model editing process.
Table 1: Full results of long sequence forecasting experiments evaluated on diverse datasets, where bold indicates the best performance and underlined indicates the second best performance. Baseline results are sourced from Woo et al. (2024a).
Zero-shot Full-shot
TimeDiT MoiraiSmall MoiraiBase MoiraiLarge iTransformer TimesNet PatchTST Crossformer TiDE DLinear SCINet FEDformer
ETTh1 (MSE) 96 0.325 0.375 0.384 0.380 0.386 0.384 0.414 0.423 0.479 0.386 0.654 0.376
192 0.347 0.399 0.425 0.440 0.441 0.436 0.460 0.471 0.525 0.437 0.719 0.420
336 0.347 0.412 0.456 0.514 0.487 0.491 0.501 0.570 0.565 0.481 0.778 0.459
720 0.404 0.413 0.470 0.705 0.503 0.521 0.500 0.653 0.594 0.519 0.836 0.506
Avg. 0.356 0.400 0.434 0.510 0.454 0.458 0.469 0.529 0.541 0.456 0.747 0.440
ETTh2 (MSE) 96 0.257 0.281 0.277 0.287 0.297 0.340 0.302 0.745 0.400 0.333 0.707 0.358
192 0.316 0.340 0.340 0.347 0.380 0.402 0.388 0.877 0.528 0.477 0.860 0.429
336 0.341 0.362 0.371 0.377 0.428 0.452 0.426 1.043 0.643 0.594 1.000 0.496
720 0.447 0.380 0.394 0.404 0.427 0.462 0.431 1.104 0.874 0.831 1.249 0.463
Avg. 0.340 0.341 0.345 0.354 0.383 0.414 0.387 0.942 0.611 0.559 0.954 0.437
ETTh1 (MAE) 96 0.386 0.402 0.402 0.398 0.405 0.402 0.419 0.448 0.464 0.400 0.599 0.419
192 0.390 0.419 0.429 0.434 0.436 0.429 0.445 0.474 0.492 0.432 0.631 0.448
336 0.396 0.429 0.450 0.474 0.458 0.469 0.466 0.546 0.515 0.459 0.659 0.465
720 0.440 0.444 0.473 0.568 0.491 0.500 0.488 0.621 0.558 0.516 0.699 0.507
Avg. 0.403 0.424 0.438 0.469 0.448 0.450 0.455 0.522 0.507 0.452 0.647 0.46
ETTh2 (MAE) 96 0.331 0.334 0.327 0.325 0.349 0.374 0.348 0.584 0.440 0.387 0.621 0.397
192 0.380 0.373 0.374 0.367 0.400 0.414 0.400 0.656 0.509 0.476 0.689 0.439
336 0.418 0.393 0.401 0.393 0.432 0.541 0.433 0.731 0.571 0.541 0.744 0.487
720 0.441 0.416 0.426 0.421 0.445 0.657 0.446 0.763 0.679 0.657 0.838 0.474
Avg. 0.393 0.379 0.382 0.376 0.407 0.497 0.407 0.684 0.550 0.515 0.723 0.449
Table 2: Forecasting results on CRPS_sum for both zero-shot and full-shot settings.
Setting Dataset TimeDiT (ZS) TEMPO MoiraiSmall MoiraiBase MoiraiLarge LagLLaMA TimeMixer TimeLLM Timer
Zero Shot Solar 0.424 0.581 0.884 0.948 1.042 0.690 0.999 0.997 0.101
Electricity 0.030 0.081 0.079 0.072 0.039 0.065 0.302 0.303 0.301
Traffic 0.351 0.147 0.215 0.191 0.111 0.275 0.403 0.368 0.384
Taxi 0.392 0.400 0.463 0.428 0.597 0.620 0.785 0.782 0.788
Exchange 0.019 0.030 0.007 0.012 0.011 0.024 0.079 0.076 0.072
Setting Dataset TimeDiT (FS) DLinear PatchTST Latent ODE GPT4TS TransMAF TimeGrad CSDI Diffusion-TS
Full Shot Solar 0.278 0.432 0.457 0.445 0.467 0.301 0.287 0.298 0.286
Electricity 0.005 0.033 0.037 0.140 0.033 0.021 0.021 0.017 0.019
Traffic 0.019 0.070 0.405 0.095 0.069 0.056 0.044 0.020 0.097
Taxi 0.123 0.177 0.190 0.181 0.187 0.179 0.114 0.123 0.303
Exchange 0.005 0.011 0.026 0.013 0.013 0.005 0.006 0.007 0.009